# Large Language Models as Realistic Microservice Trace Generators

### v2

@@ Abstract
Workload traces are essential to understand complex computer systems’ behavior and manage processing and memory resources. Since real-world traces are hard to obtain, synthetic trace generation is a promising alternative. This paper proposes a first-of-a-kind approach that relies on training a large language model (LLM) to generate synthetic workload traces, specifically microservice call graphs. To capture complex and arbitrary hierarchical structures and implicit constraints in such traces, we show how to fine-tune LLMs to generate recursively, making call graph generation a sequence of easier steps. To further enforce learning constraints in traces and generate uncommon situations, we argue for applying additional instruction tuning steps to align our model with the desired trace features. Our evaluation results show that we can generate diverse realistic traces under various conditions and outperform existing methods in accuracy and validity. We demonstrate that our synthetically generated traces can effectively replace real data to optimize important microservice management tasks. Additionally, our model adapts to downstream trace-related tasks, such as predicting key trace features and infilling missing data.
@@
Donghyun Kim, Sriram Ravula, Taemin Ha, Alexandros G. Dimakis, Daehyeok Kim, Aditya Akella University of Texas at Austin {donghyun,sriram.ravula,taemin.ha,dimakis}@utexas.edu, {daehyeok,akella}@cs.utexas.edu
@@ 1 Introduction
In this paper, we show how to use large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023) to generate synthetic workload traces. LLMs have been successfully adapted beyond natural language to domains like protein sequences (Shen et al., 2024), code (Roziere et al., 2023), and tabular data (Borisov et al., 2023). In addition, LLMs can produce outputs well-aligned with user inputs through fine-tuning (Ouyang et al., 2022; Wei et al., 2021) and generalize to new prompts at inference (Chung et al., 2022; Sanh et al., 2021). Thus, we posit that LLMs have the potential to generate synthetic traces that accurately capture real-world systems trace structures while adhering to user specifications.
@@
Our approach adapts general-purpose LLMs to generate synthetic system traces, particularly focusing on microservice call graphs. Microservices are the de facto approach to designing modern cloud-based applications and power many popular services today (Uber, ; Netflix, ; Amazon, ). Microservice call graphs trace the execution of user requests through such applications. These call graphs form a type of trace with a rich directed acyclic graph (DAG) structure. We represent these graphs in a text-based format suitable for LLMs, enabling their use in trace generation. Crucially, one of our key innovations is to generate call graphs with structural constraints by recursively generating subgraphs, or layers. This approach allows the model to break down the complex task of reasoning about hierarchical graph structures and complex constraints into multiple simpler tasks. To further enhance the model’s ability to follow structural constraints and to meet user-requested attributes, we employ instruction tuning. During this phase, the model learns to explicitly generate a series of intermediate instructions between recursive layer generation steps, performing arithmetic and logical checks to ensure strict adherence to the desired structure.
We validate our approach by fine-tuning Llama-2 7B (Touvron et al., 2023) on microservice trace data and conducting extensive evaluations. Results show that recursive generation with intermediate instructions significantly improves the model’s ability to generate valid outputs for complex call graphs. Compared to generative and probabilistic models, our synthetic traces better align with real trace distributions. Furthermore, we demonstrate that synthetic traces effectively replace real data for training microservice management tasks and that our fine-tuned model excels Llama-3.1 405B (Dubey et al., 2024) in trace feature prediction. We release our codes in https://github.com/ldos-project/TraceLLM.
@@ 3 Training LLMs for Microservice Traces
Our goal is to train a generative model for microservice call graph traces, enabling end-users to simulate diverse scenarios like rare microservice invocations that exhibit high response times. To achieve this, we condition the model’s output on user-specified attributes, including the invoked application, the number of microservice communications (i.e., graph edges), and overall application latency. Given the limitations of existing trace generation methods, we leverage LLMs. We initialize our model from LLMs pre-trained on large text datasets, as these models have proven effective when adapted to specialized domains such as proteins (Shen et al., 2024) and code (Roziere et al., 2023). Moreover, LLMs support flexible conditioning mechanisms, including natural language prompts (Ouyang et al., 2022) and structured input sequences (Borisov et al., 2023).
This section describes our approach to training an LLM for microservice call graph generation in two stages. First, we pre-train the model to capture the interactions in real-world call graph data and introduce a recursive subgraph generation to enhance its ability to generate large, structured graphs. Second, we fine-tune the model with instruction tuning, enabling it to generate call graphs with user-specified attributes and ensuring constraint adherence through natural language reasoning.
@@ 3.1.2 Recursive Generation
Formally, for an encoded call graph $\mathbf{t}=[\mathbf{t}_{1},\mathbf{t}_{2},\dots,\mathbf{t}_{n}]$, we partition the edges $\mathbf{t}_{i}$ into a sequence of layers $[\mathbf{t}^{1},\mathbf{t}^{2},\dots,\mathbf{t}^{l}]$, where $l\leq n$. Each layer consists of a sequence of edges that share the same parent (i.e., source) node, ensuring that no edge appears in multiple layers. For call graph conditions $\mathbf{c}$ that describe $\mathbf{t}$, we introduce layer conditions $\mathbf{c}^{j}$, $j\in\{1,2,\dots,l+1\}$. Layer condition $\mathbf{c}^{j}$ encodes the attributes of the remaining portion of the call graph after the sequence of layers $[\mathbf{t}^{1},\mathbf{t}^{2},\dots,\mathbf{t}^{j-1}]$ has been generated, and we define $\mathbf{c}^{1}\coloneqq\mathbf{c}$ and $\mathbf{c}^{l+1}\coloneqq\emptyset$. We decompose the conditional call graph distribution as a chain of conditional layer distributions: $p(\mathbf{t}|\mathbf{c})=\prod_{k=1}^{l}p(\mathbf{c}^{k+1},\mathbf{t}^{k}|\mathbf{c}^{k})$. In other words, the model predicts call graphs from user prompts iteratively layer-by-layer. For layer $k$ the model takes $\mathbf{c}^{k}$ as input and produces the sequence of edges $\mathbf{t}^{k}$ followed by the conditions $\mathbf{c}^{k+1}$ of the next layer. The process continues recursively, using $\mathbf{c}^{k+1}$ to predict the next layer, $k+1$. Figure 2 illustrates an example of this recursive generation.
@@ 4 Evaluation
We thoroughly demonstrate the effectiveness of our method in two major aspects: (1) synthetic trace quality in terms of structural validity (§ 4.1), distribution similarity (§ 4.2), and usefulness to train and evaluate machine learning-driven microservice management tasks (§ 4.3), and (2) benefits from our use of LLMs in terms of instruction-following capabilities (§ 4.4) and trace-related downstream task performance (§ 4.5).
@@ 4.1 Structured Reasoning Accuracy
Results.  3(a) and 3(b) present the accuracy of generated call graphs across varying numbers of edges and depths. Generally, as complexity increases (i.e., more edges or greater depth), the baseline model’s accuracy decreases significantly—dropping below 25% for edges $>$ 15 and nearing zero for depths $>$ 4. In contrast, the recursive generation model maintains higher accuracies, by approximately 30% and 35%, respectively. This improved performance is attributed to the model breaking down complex generation tasks into simpler, more manageable sub-tasks.
@@ 4.2 Similarity of Real and Synthetic Traces
4(b) illustrates the similarity for varying $K$ values, from 10 to 500. Our method achieves over 90% accuracy for $K\leq 50$ and 65% at $K$=500, demonstrating robust performance. The model trained with the GReaT method also shows robust performance, but slightly worse performance with larger $K$ values. We believe that the performance gap results from the lack of capability to generate complex structures (§ 4.1), which impacts trace distribution. In contrast, the probabilistic model starts at 59% accuracy for $K$=10 and declines to 23% at $K$=500, showcasing our method’s capability to capture and replicate heavy-hitter dynamics.
@@ 4.3 Synthetic Data as Training data for ML-Driven Microservice Management
Critical Component Extraction.  FIRM (Qiu et al., 2020) predicts critical components (microservices likely to violate service level objectives (SLO)) using support vector machines (SVMs) trained on latency-related features from call graphs. For our evaluation, we train SVMs to detect critical components using two popular applications (apps A and B) from our trace dataset. For each application, we randomly sample call graphs and train two SVMs: one with real data and one with synthetic data generated by our fine-tuned model. As shown in Figure 5, SVMs trained on synthetic data achieve near-identical accuracy to those trained on real data, with a difference of less than 1.5 percentage points. In contrast, SVMs trained on synthetic traces from baselines show a performance gap of 6 to 81 percentage points.
Anomaly Detection.  For operators to efficiently diagnose system failures, anomaly detection models predict whether microservice call graphs include anomalous characteristics like irregular graph structure or time. We assess our synthetic data quality using TraceVAE (Xie et al., 2023), a variational autoencoder (VAE) model that detects anomalous microservices in terms of time consumption. We train TraceVAE models using real and synthetic trace data, similar to our previous experiment. Figure 5 shows that models trained on synthetic traces from our method yield results comparable to those trained on real data, as measured by ROC AUC.
@@ 5 Conclusion
This paper introduces a training method for adapting pre-trained LLMs to generate microservice trace graphs using recursive call graph generation and instruction tuning. Our approach outperforms baselines in producing accurate, valid call graphs with improved distributional similarity to real traces. We demonstrate that synthetic traces can effectively replace real data for training microservice management tasks, such as critical component extraction and anomaly detection. Additionally, instruction tuning enhances graph generation based on user-specified features, enabling applications in prediction and data infilling. While we focus on microservice call graphs, our method broadly applies to other system traces with similar structures.
6 Limitations
This section discusses a few limitations of our work and potential approaches to overcome them.
@@ 6 Limitations
The recursive method improves the accuracy of call graph generation compared to generating the entire trace at once, but a key drawback is that previously generated edges are discarded, as only the conditioning information from the prior layer is passed to the next layer generation steps. Although dropping previously generated edges has little impact on the output in microservice call graph generation, where direct neighbors exert the most influence (Zhang et al., 2024), incorporating past information, such as prior layers or a time series of call graph traces, could enhance the capture of longer-range dependencies and temporal patterns. However, efficiently compressing historical trace information while preserving critical details remains an open challenge. In future work, we will consider this direction to compress long-range traces and generate synthetic traces conditioned on the compressed traces.
@@
Furthermore, our method uses manually constructed instruction templates, which may lead to suboptimal generation quality, as we are not using the full potential of language models pre-trained with trillions of tokens (Touvron et al., 2023). Following the methods of prior work Liu et al. (2024); Gunasekar et al. (2023); Li et al. (2024), we believe that diversifying instructions using LLM-generated output is a potential method to improve the ability of LLMs to follow user intentions. However, naively guiding LLMs to generate instructions for trace generation may result in instructions that lack useful characteristics for downstream tasks. In future work, we plan to integrate domain-specific knowledge of traces to improve the usefulness and diversity of instructions generated by LLMs.
Lastly, we focused on generating microservice call graphs in this paper, but other system traces, such as operating system (OS) call graphs, share a similar hierarchical structure. The primary differences in OS call graphs lie in their greater depth and the increased diversity of node and edge types. Evaluating whether our approach can be effectively applied to such traces remains an open question.
@@ B.1 Training Setup
We train all models with 4$\times$ A100 80GB GPUs in our cluster with the hyperparameters described in Table 1. We apply LoRA ((Hu et al., 2022)) adapters to query and key projection matrices of attention layers with $rank=8$, $alpha=16$, and $dropout=0.1$. For the downstream task training in § 4.5, we freeze the backbone model and only train the last classification layer for the prediction task. For the infilling downstream task, we use LoRA adapters with the same configuration as mentioned earlier. During inference, we use a temperature of 0.8 and top-K of 50 for trace generation, unless otherwise specified. We use models (Llama 2 and 3.1) under appropriate community license.
@@ B.3 Training Data Examples
From the call graph traces, we create text-based representations of call graphs as described in Section 3.1.1. First of all, Figure 9 is a training data example of converting a call graph into a tabular data format, which is the baseline in § 4.1. At the beginning, we provide high-level information about the call graph, including the service ID, number of edges, and graph depth. Each line inside the <edges> block corresponds to a call in a call graph. 6 fields exist for each call including the edge ID, source/destination microservices, communication type, and communication start/finish time.
Figure 10 shows an example training data sample for recursive generation as described in Section 3.1.2. Each sample consists of a sequence of layers, where each layer includes the edges and the conditions for the next layers. At the beginning of each layer, we provide high-level information to explain connections with the previous layers (e.g., start_node, caller), structure in the call graph (e.g., remaining_depth, num_edges, start_edge_id), and time-related information (e.g., latency, start_communication_at). Note that the number of fields in each edge is reduced from 6 to 5 since the edges share the same start node. Also, the edge ID field is an integer, not a dot-decimal number. For each next layer, the condition is described in each <subgraph> block starting with the edge ID to be extended.
@@
As described in Section 3.1.1, we drop each call graph attribute randomly with probability $p_{drop}$. We set $p_{drop}$ to 0.9 for all attributes except for the service ID field, which is always kept ($p_{drop}=1$), to ensure minimal conditioning.

### v3

@@ Abstract
Workload traces are essential to understand complex computer systems’ behavior and manage processing and memory resources. Since real-world traces are hard to obtain, synthetic trace generation is a promising alternative. This paper proposes a first-of-a-kind approach that relies on training a large language model (LLM) to generate synthetic workload traces, specifically microservice call graphs. To capture complex and arbitrary hierarchical structures and implicit constraints in such traces, we propose to train LLMs to generate recursively, making call graph generation a sequence of more manageable steps. To further enforce learning constraints on the traces and generate uncommon situations, we apply additional instruction tuning steps to align our model with the desired trace features. With this method, we train TraceLLM, an LLM for microservice trace generation, and demonstrate that it produces diverse, realistic traces under varied conditions, outperforming existing approaches in both accuracy and validity. The synthetically generated traces can effectively replace real data to optimize important microservice management tasks. Additionally, TraceLLM adapts to downstream trace-related tasks, such as predicting key trace features and infilling missing data.
@@ 1 Introduction
In this paper, we propose TraceLLM that adapts pre-trained large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023) to generate synthetic workload traces. LLMs have been successfully adapted beyond natural language to domains like protein sequences (Shen et al., 2024), code (Roziere et al., 2023), and tabular data (Borisov et al., 2023). In addition, LLMs can produce outputs well-aligned with user inputs through fine-tuning (Ouyang et al., 2022; Wei et al., 2021) and generalize to new prompts at inference (Chung et al., 2022; Sanh et al., 2021). Thus, we posit that LLMs have the potential to generate synthetic traces that accurately capture real-world systems trace structures while adhering to user specifications.
@@
In this paper, we design TraceLLM to address these challenges, with a particular focus on generating microservice call graphs. Microservices are the de facto approach to designing modern cloud-based applications and power many popular services today (Uber, ; Netflix, ; Amazon, ). Microservice call graphs trace the execution of user requests through such applications. These call graphs form a type of trace with a rich directed acyclic graph (DAG) structure. TraceLLM represents these graphs in a text-based format suitable for LLMs, enabling their use in trace generation. Crucially, one of our key innovations is to generate call graphs with structural constraints by recursively generating subgraphs, or layers. This approach allows the model to break down the complex task of reasoning about hierarchical graph structures and complex constraints into multiple simpler tasks. To further enhance the model’s ability to follow structural constraints and to meet user-requested attributes, we employ instruction tuning. During this phase, the model learns to explicitly generate a series of intermediate instructions between recursive layer generation steps, performing arithmetic and logical checks to ensure strict adherence to the desired structure.
We implement TraceLLM by training Llama-2 7B (Touvron et al., 2023) on microservice trace data and evaluate it through extensive evaluations. Results show that recursive generation with intermediate instructions significantly improves the model’s ability to generate valid outputs for complex call graphs. Compared to generative and probabilistic models, our synthetic traces better align with real trace distributions. Furthermore, we demonstrate that synthetic traces effectively replace real data for training microservice management tasks and that our fine-tuned model excels Llama-3.1 405B (Dubey et al., 2024) in trace feature prediction. We release our codes in https://github.com/ldos-project/TraceLLM.
@@ 3 Training LLMs for Microservice Traces
Our goal is to train TraceLLM, an LLM for microservice call graph traces, enabling end-users to simulate diverse scenarios, such as rare microservice invocations that exhibit high response times. To achieve this, we condition the model’s output on user-specified attributes, including the invoked application, the number of microservice communications (i.e., graph edges), and overall application latency. Given the limitations of existing trace generation methods, we leverage LLMs. We initialize our model from LLMs trained on large text datasets, as these models have proven effective when adapted to specialized domains such as proteins (Shen et al., 2024) and code (Roziere et al., 2023). Moreover, LLMs support flexible conditioning mechanisms, including natural language prompts (Ouyang et al., 2022) and structured input sequences (Borisov et al., 2023).
This section presents our two-stage approach for training TraceLLM to generate microservice call graphs. First, we extend LLMs with additional pre-training on a large corpus of microservice call graph traces, allowing the model to capture interaction patterns in real-world call graphs. We also introduce recursive subgraph generation to improve its ability to produce large, structured graphs. Second, we instruction fine-tune the model, enabling it to generate call graphs with user-specified attributes and ensuring constraint adherence through natural language reasoning.
@@ 3.1.2 Recursive Generation
Formally, for an encoded call graph $\mathbf{t}=[\mathbf{t}_{1},\mathbf{t}_{2},\dots,\mathbf{t}_{n}]$, we partition the edges $\mathbf{t}_{i}$ into a sequence of layers $\mathbf{t}^{1},\mathbf{t}^{2},\dots,\mathbf{t}^{l}]$, where $l\leq n$. Each layer consists of a sequence of edges that share the same parent (i.e., source) node, ensuring that no edge appears in multiple layers. For call graph conditions $\mathbf{c}$ that describe $\mathbf{t}$, we introduce layer conditions $\mathbf{c}^{j}$, $j\in\{1,2,\dots,l+1\}$. Layer condition $\mathbf{c}^{j}$ encodes the attributes of the remaining portion of the call graph after the sequence of layers $\mathbf{t}^{1},\mathbf{t}^{2},\dots,\mathbf{t}^{j-1}]$ has been generated, and we define $\mathbf{c}^{1}\coloneqq\mathbf{c}$ and $\mathbf{c}^{l+1}\coloneqq\emptyset$. We decompose the conditional call graph distribution as a chain of conditional layer distributions: $p(\mathbf{t}|\mathbf{c})=\prod_{k=1}^{l}p(\mathbf{c}^{k+1},\mathbf{t}^{k}|\mathbf{c}^{k})$. In other words, the model predicts call graphs from user prompts iteratively layer-by-layer. For layer $k$ the model takes $\mathbf{c}^{k}$ as input and produces the sequence of edges $\mathbf{t}^{k}$ followed by the conditions $\mathbf{c}^{k+1}$ of the next layer. The process continues recursively, using $\mathbf{c}^{k+1}$ to predict the next layer, $k+1$. Figure 2 illustrates an example of this recursive generation.
@@ 4 Evaluation
We thoroughly demonstrate the effectiveness of TraceLLM in two major aspects: (1) synthetic trace quality in terms of structural validity (§ 4.1), distribution similarity (§ 4.2), and usefulness to train and evaluate machine learning-driven microservice management tasks (§ 4.3), and (2) benefits from our use of LLMs in terms of instruction-following capabilities (§ 4.4) and trace-related downstream task performance (§ 4.5).
@@ 4.1 Structured Reasoning Accuracy
Results.  3(a) and 3(b) present the accuracy of generated call graphs across varying numbers of edges and depths. Generally, as complexity increases (i.e., more edges or greater depth), the baseline model’s accuracy decreases significantly—dropping below 25% for edges >> 15 and nearing zero for depths >> 4. In contrast, the recursive generation model maintains higher accuracies, by approximately 30% and 35%, respectively. This improved performance is attributed to the model breaking down complex generation tasks into simpler, more manageable sub-tasks.
@@ 4.2 Similarity of Real and Synthetic Traces
4(b) illustrates the similarity for varying $K$ values, from 10 to 500. Our method achieves over 90% accuracy for $K\leq 50$ and 65% at KK=500, demonstrating robust performance. The model trained with the GReaT method also shows robust performance, but slightly worse performance with larger $K$ values. We believe that the performance gap results from the lack of capability to generate complex structures (§ 4.1), which impacts trace distribution. In contrast, the probabilistic model starts at 59% accuracy for KK=10 and declines to 23% at KK=500, showcasing our method’s capability to capture and replicate heavy-hitter dynamics.
@@ 4.3 Synthetic Data as Training data for ML-Driven Microservice Management
Critical Component Extraction.  FIRM (Qiu et al., 2020) predicts critical components (microservices likely to violate service level objectives (SLO)) using support vector machines (SVMs) trained on latency-related features from call graphs. For our evaluation, we train SVMs to detect critical components using two popular applications (apps A and B) from our trace dataset. For each application, we randomly sample call graphs and train two SVMs: one with real data and one with synthetic data generated by our fine-tuned model. As shown in Figure˜5, SVMs trained on synthetic data achieve near-identical accuracy to those trained on real data, with a difference of less than 1.5 percentage points. In contrast, SVMs trained on synthetic traces from baselines show a performance gap of 6 to 81 percentage points.
Anomaly Detection.  For operators to efficiently diagnose system failures, anomaly detection models predict whether microservice call graphs include anomalous characteristics like irregular graph structure or time. We assess our synthetic data quality using TraceVAE (Xie et al., 2023), a variational autoencoder (VAE) model that detects anomalous microservices in terms of time consumption. We train TraceVAE models using real and synthetic trace data, similar to our previous experiment. Figure˜5 shows that models trained on synthetic traces from our method yield results comparable to those trained on real data, as measured by ROC AUC.
@@ 5 Conclusion
This paper introduces TraceLLM, a training method for adapting LLMs to generate microservice trace graphs using recursive call graph generation and instruction tuning. Our approach outperforms baselines in producing accurate, valid call graphs with improved distributional similarity to real traces. We demonstrate that synthetic traces can effectively replace real data for training microservice management tasks, such as critical component extraction and anomaly detection. Additionally, instruction tuning enhances graph generation based on user-specified features, enabling applications in prediction and data infilling. While we focus on microservice call graphs, our method broadly applies to other system traces with similar structures.
@@
Limitations.  The recursive method improves the accuracy of call graph generation compared to generating the entire trace at once, but a key drawback is that previously generated edges are discarded, as only the conditioning information from the prior layer is passed to the next layer generation steps. Although dropping previously generated edges has little impact on the output in microservice call graph generation, where direct neighbors exert the most influence (Zhang et al., 2024), incorporating past information, such as prior layers or a time series of call graph traces, could enhance the capture of longer-range dependencies and temporal patterns. However, efficiently compressing historical trace information while preserving critical details remains an open challenge. In future work, we will consider this direction to compress long-range traces and generate synthetic traces conditioned on the compressed traces.
@@
Furthermore, our method uses manually constructed instruction templates, which may lead to suboptimal generation quality, as we are not using the full potential of language models pre-trained with trillions of tokens (Touvron et al., 2023). Following the methods of prior work (Liu et al., 2024; Gunasekar et al., 2023; Li et al., 2024), we believe that diversifying instructions using LLM-generated output is a potential method to improve the ability of LLMs to follow user intentions. However, naively guiding LLMs to generate instructions for trace generation may result in instructions that lack useful characteristics for downstream tasks. In future work, we plan to integrate domain-specific knowledge of traces to improve the usefulness and diversity of instructions generated by LLMs.
Lastly, we focused on generating microservice call graphs in this paper, but other system traces, such as operating system (OS) call graphs, share a similar hierarchical structure. The primary differences in OS call graphs lie in their greater depth and the increased diversity of node and edge types. We present relevant experimental results in § D.7, using cluster batch job traces. We believe that extending TraceLLM’s capabilities to diverse types of system traces is important, since TraceLLM can function not only as a synthetic trace generator but also as a world model (Team et al., 2025) that offers realistic feedback for developing more sophisticated agents capable of reasoning over various computer system traces.
@@
Acknowledgments.  This material is based upon work supported by the U.S. National Science Foundation (NSF) under Grant Number 2326576. We acknowledge the use of AI assistants to enhance writing clarity.
@@ B.1 Training Setup
[…]
@@ B.3 Training Data Examples
From the call graph traces, we create text-based representations of call graphs as described in Section˜3.1.1. First of all, Figure 9 is a training data example of converting a call graph into a tabular data format, which is the baseline in § 4.1. At the beginning, we provide high-level information about the call graph, including the service ID, number of edges, and graph depth. Each line inside the <edges> block corresponds to a call in a call graph. 6 fields exist for each call including the edge ID, source/destination microservices, communication type, and communication start/finish time.
Figure 10 shows an example training data sample for recursive generation as described in Section˜3.1.2. Each sample consists of a sequence of layers, where each layer includes the edges and the conditions for the next layers. At the beginning of each layer, we provide high-level information to explain connections with the previous layers (e.g., start_node, caller), structure in the call graph (e.g., remaining_depth, num_edges, start_edge_id), and time-related information (e.g., latency, start_communication_at). Note that the number of fields in each edge is reduced from 6 to 5 since the edges share the same start node. Also, the edge ID field is an integer, not a dot-decimal number. For each next layer, the condition is described in each <subgraph> block starting with the edge ID to be extended.
@@
As described in Section˜3.1.1, we drop each call graph attribute randomly with probability $p_{drop}$. We set $p_{drop}$ to 0.9 for all attributes except for the service ID field, which is always kept ($p_{drop}=1$), to ensure minimal conditioning.